A rectangular parking lot has a length that is 3 yards greater than the width. The area of the parking lot is 700 square yards. Find the length and the width. Use the formula, area=length*width. Can you please show all the steps so that I can see how you are doing this problem? Thank you!!
Answer by checkley77(12844)

L=3W
700=L*W
700=3W*W
3W^2=700
W^2=700/3
W^2=233.33
W=SQRT233.33
W=15.275 ANS.
L=3*15.275
L=45.826 ANS.
PROOF
700=45.826*15.275
700=699.988